# “Give Me an Example Like This”: Episodic Active Reinforcement Learning from Demonstrations

### v2

@@ III-C Trajectory-Based Uncertainty Measurement
III-D Episodic Active Reinforcement Learning from Demonstration Query (EARLY)
Utilizing the trajectory-based uncertainty measurement in Section III-C and the trajectory-based feature space introduced in Section III-A, we construct an active query strategy for episodic expert demonstrations to solve the problems of when to query and what to query.
During each training iteration, we first sample an initial state $s_{0}^{i}$, obtain an episodic roll-out trajectory $\xi_{\pi}^{i}$ by the current agent policy $\pi$, and calculate its corresponding feature value $\varphi_{i}$. To evaluate how the learning agent is uncertain for this feature point, we estimate the uncertainty $u_{i}$ of the obtained feature point $\varphi_{i}$ as the agent uncertainty along this generated roll-out trajectory $\xi_{i}^{\pi}$ (i.e., $u_{i}=u(\xi_{i}^{\pi})$). Both the sampled feature point $\varphi_{i}$ and its corresponding uncertainty estimation $u_{i}$ will be stored in shifting recent histories, one for feature points and one for uncertainty values. After the shifting recent history grows to its full length $N_{h}$, an adaptive uncertainty threshold will be determined via a ratio threshold $r_{query}\in[0,1]$ as in  [7]. Whenever the current uncertainty value $u_{i}$ is among the top $r_{query}$ of the shifting recent history of uncertainty and the demonstration query budget $N_{d}$ has not been used up, the learning agent will decide to make a query for one episodic expert demonstration.
Different from  [7], we choose to query the most uncertain feature point $\varphi_{query}$ in the shifting recent history and ask for an episodic expert demonstration $\xi_{\pi^{demo}}^{k}$, whose feature value is expected to be the same as the queried feature point $\varphi_{query}$. Both the learning policy roll-out $\xi_{\pi}^{i}$ and the expert episodic demonstration $\xi_{\pi_{demo}}^{k}$ will be added to the reply buffer $D$ to update agent policy using SAC as the underlying RL algorithm. We summarize the pseudo-code in Algorithm 1.
@@ III-D Episodic Active Reinforcement Learning from Demonstration Query (EARLY)
13:     calculate uncertainty $u_{i}\leftarrow M(\xi_{\pi}^{i},Q_{\theta_{k}},V_{\psi},V_{\bar{\psi}},\pi_{\phi})$
@@
16:         ordered history $H_{u}^{dsc}\leftarrow DescOrder(H_{u})$
@@
20:              get an expert demo $\xi_{\pi^{demo}}^{k}$ of feature $\varphi_{query}$
@@ IV-A Task Environments
Figure 2: Task environments for three simulated navigation tasks of scaling difficulties.
IV-B Baselines
@@
  1. 1.
    DDPG-LfD: a popular method for reinforcement learning from demonstrations [2]. The agent learns in a conventional “demo-then-training” manner, where episodic expert demonstrations are first randomly collected and added to the reply buffer before the learning agent starts to update its control policy using DDPG.
  2. 2.
    I-ARLD: a state-of-the-art method that learns in a “demo-while-training” manner [8]. It switches control from the learning agent to the expert demonstrator during the agent roll-outs, resets the environment to previous moments, and only queries isolated state-action pairs for the next few steps before switching control back to the learning agent.
  3. 3.
    GAIL: a classic imitation learning algorithm that also learns in a “demo-then-training” manner [28].
  4. 4.
    BC: one of the most common imitation learning algorithms that directly treats policy training as a conventional supervised learning problem [29].
@@
Figure 3: Results of the experiments with simulated-oracle demonstrators. The shaded areas represent the $95\%$ confidence intervals.
For our method, we chose the ratio threshold $r_{query}$ as $0.35,0.4$, and $0.55$ for three navigation tasks respectively, and set the maximum length of recent explored history $N_{h}$ as $20$. For the underlying SAC algorithm, we followed the same settings of neural network structures, hyperparameters, and the optimizer as in  [26]. For DDPG-LfD and I-ARLD, we reproduced them according to their original papers with the default parameters. For GAIL and BC, we implemented them using the open-source library [30] for stable implementation. For all baselines, we trained the policy with $1\times 10^{5}$ environment steps (i.e., $i_{max}$) for all three tasks respectively.
Additionally, we did not find performance improvement by using Prioritized Experience Replay (PER)  [15] for the reply buffer. Instead, we maintained two separate reply buffers for current policy roll-outs and expert demonstrations. To guarantee the expert demonstrations can be stably sampled, we sampled the same amount of roll-outs from expert demonstrations as those from the agent policy to comprise each sampling batch. All expert demonstrations will be stored in the corresponding reply buffer through the whole learning process, while the earliest agent roll-out will be removed from the reply buffer for the agent policy once it exceeds the buffer capacity.
IV-C Experiments with Oracle-Simulated Demonstrators
We first conducted experiments using oracle-simulated demonstrators to evaluate the learning performance of our method. We used RRT*  [31], a state-of-the-art path planning algorithm, as the oracle to provide episodic demonstrations upon receiving feature-oriented queries from the learning agent. Since we chose the initial state as the feature $\varphi_{i}$ of a given episodic roll-out trajectory $\xi_{\pi}^{i}$, whenever a feature query $\varphi_{query}$ (i.e., $s_{0}^{query}$) was generated, we intuitively used the RRT* algorithm to obtain an episodic expert roll-out trajectory that starts from $s_{0}^{query}$ and arrives at the destination. For the baselines that learn in a “demo-then-training” manner (i.e., DDPG-LfD, GAIL, and BC), we uniformly sampled from the initial state space to select the initial states of the expert demonstrations. To keep data collection labor aligned with a reasonable amount for real human demonstrators, we only allowed the learning agent to query $60$ episodic expert demonstrations (i.e., $N_{d}=60$) for each baseline method (or of an equal amount of total steps for I-ARLD).
@@ IV-C Experiments with Oracle-Simulated Demonstrators
Figure 4: Results of experiments with real human demonstrators.
IV-D Pilot User Study with Human Expert Demonstrators
To investigate the efficacy of our algorithm and its user experience for real human users, we conducted a pilot user study with $18$ human participants ($9$ male, $8$ female, and $1$ other; $12$ aged between $18-29$ and $6$ aged between $30-39$; $11$ of some experience of machine learning and $7$ of extensive experience). We recruited them from campus via poster advertisement following the ethical guidelines provided by our faculty’s research ethics board. We obtained their consent for experiments and data collection before the experiments began and compensated for their participation with a €$10$ gift card.
@@ IV-D Pilot User Study with Human Expert Demonstrators
V Results and Discussion
V-A Experiments with Oracle Experts
@@
V-B Experiments with Human Experts
V-B1 Learning Performance
@@
Figure 5: Distribution of provided demonstrations from one of the human participants using different baseline methods.

### v3

@@ 3.3. Trajectory-Based Uncertainty Measurement
3.4. Episodic Active Reinforcement Learning from Demonstration Query (EARLY)
Utilizing the trajectory-based uncertainty measurement in Section 3.3 and the trajectory-based feature space introduced in Section 3.1, we construct an active query strategy for episodic expert demonstrations to solve the problems of when to query and what to query.
During each training iteration, we first sample an initial state $s_{0}^{i}$, obtain an episodic roll-out trajectory $\xi_{\pi}^{i}$ by the current agent policy $\pi$, and calculate its corresponding feature value $\varphi_{i}$. To evaluate how the learning agent is uncertain for this feature point, we estimate the uncertainty $u_{i}$ of the obtained feature point $\varphi_{i}$ as the agent uncertainty along this generated roll-out trajectory $\xi_{i}^{\pi}$ (i.e., $u_{i}=u(\xi_{i}^{\pi})$). Both the sampled feature point $\varphi_{i}$ and its corresponding uncertainty estimation $u_{i}$ will be stored in shifting recent histories, one for feature points and one for uncertainty values. After the shifting recent history grows to its full length $N_{h}$, an adaptive uncertainty threshold will be determined via a ratio threshold $r_{query}\in[0,1]$ as in  (Chen et al., 2020). Whenever the current uncertainty value $u_{i}$ is among the top $r_{query}$ of the shifting recent history of uncertainty and the demonstration query budget $N_{d}$ has not been used up, the learning agent will decide to make a query for one episodic expert demonstration.
Different from  (Chen et al., 2020), we choose to query the most uncertain feature point $\varphi_{query}$ in the shifting recent history and ask for an episodic expert demonstration $\xi_{\pi^{demo}}^{k}$, whose feature value is expected to be the same as the queried feature point $\varphi_{query}$. Both the learning policy roll-out $\xi_{\pi}^{i}$ and the expert episodic demonstration $\xi_{\pi_{demo}}^{k}$ will be added to the reply buffer $D$ to update agent policy using SAC as the underlying RL algorithm. We summarize the pseudo-code in Algorithm 1.
@@ 3.4. Episodic Active Reinforcement Learning from Demonstration Query (EARLY)
13:     calculate feature uncertainty $u_{i}\leftarrow M(\xi_{\pi}^{i},Q_{\theta_{k}},V_{\psi},V_{\bar{\psi}},\pi_{\phi})$
@@
16:         ordered uncertainty history $H_{u}^{dsc}\leftarrow DescOrder(H_{u})$
@@
20:              get an expert demo $\xi_{\pi^{demo}}^{k}$ of feature value $\varphi_{query}$
@@ 4.1. Task Environments
Figure 2. Task environments for three simulated navigation tasks of scaling difficulties.
4.2. Baselines
@@
  1. (1)
    DDPG-LfD: a popular method for reinforcement learning from demonstrations (Vecerik et al., 2017). The agent learns in a conventional “demo-then-training” manner, where episodic expert demonstrations are first randomly collected and added to the reply buffer before the learning agent starts to update its control policy using DDPG.
  2. (2)
    I-ARLD: a state-of-the-art method that learns in a “demo-while-training” manner (Chen et al., 2023). It switches control from the learning agent to the expert demonstrator during the agent roll-outs, resets the environment to previous moments, and only queries isolated state-action pairs for the next few steps before switching control back to the learning agent.
  3. (3)
    GAIL: a classic imitation learning algorithm that also learns in a “demo-then-training” manner (Ho and Ermon, 2016).
  4. (4)
    BC: one of the most common imitation learning algorithms that directly treats policy training as a conventional supervised learning problem (Pomerleau, 1988).
@@
Figure 3. Results of the experiments with simulated-oracle demonstrators. The shaded areas represent the standard deviation.
For our method, we chose the ratio threshold $r_{query}$ as $0.35,0.4$, and $0.55$ for three navigation tasks respectively, and set the maximum length of recent explored history $N_{h}$ as $20$. For the underlying SAC algorithm, we followed the same settings of neural network structures, hyperparameters, and the optimizer as in  (Haarnoja et al., 2018). For DDPG-LfD and I-ARLD, we reproduced them according to their original papers with the default parameters. For GAIL and BC, we implemented them using the open-source library (Gleave et al., 2022) for stable implementation. For all baselines, we trained the policy with $1\times 10^{5}$ environment steps for all three tasks respectively.
Additionally, we did not find performance improvement by using Prioritized Experience Replay (PER)  (Schaul et al., 2015) for the reply buffer. Instead, we maintained two separate reply buffers for current policy roll-outs and expert demonstrations. To guarantee the expert demonstrations can be stably sampled, we sampled the same amount of roll-outs from expert demonstrations as those from the agent policy to comprise each sampling batch. All expert demonstrations will be stored in the corresponding reply buffer through the whole learning process, while the earliest agent roll-out will be removed from the reply buffer for the agent policy once it exceeds the buffer capacity.
4.3. Experiments with Oracle-Simulated Demonstrators
We first conducted experiments using oracle-simulated demonstrators to evaluate the learning performance of our method. We used RRT*  (Karaman and Frazzoli, 2010), a state-of-the-art path planning algorithm, as the oracle to provide episodic demonstrations upon receiving feature-oriented queries from the learning agent. Since we chose the initial state as the feature $\varphi_{i}$ of a given episodic roll-out trajectory $\xi_{\pi}^{i}$, whenever a feature query $\varphi_{query}$ (i.e., $s_{0}^{query}$) was generated, we intuitively used the RRT* algorithm to obtain an episodic expert roll-out trajectory that starts from $s_{0}^{query}$ and arrives at the destination. For the baselines that learn in a “demo-then-training” manner (i.e., DDPG-LfD, GAIL, and BC), we uniformly sampled from the initial state space to select the initial states of the expert demonstrations. To keep data collection labor aligned with a reasonable amount for real human demonstrators, we only allowed the learning agent to query $60$ episodic expert demonstrations (i.e., $N_{d}=60$) for each baseline method (or of an equal amount of total steps for I-ARLD).
@@ 4.3. Experiments with Oracle-Simulated Demonstrators
Figure 4. Results of experiments with real human demonstrators.
4.4. Pilot User Study with Human Expert Demonstrators
To investigate the efficacy of our algorithm and its user experience for real human users, we conducted a pilot user study with $18$ human participants ($9$ male, $8$ female, and $1$ other; $12$ aged between $18-29$ and $6$ aged between $30-39$; $11$ of some experience of machine learning and $7$ of extensive experience). We recruited them from campus via poster advertisement with the approval of our faculty research ethics board. We obtained their consent for experiments and data collection before the experiments began and compensated for their participation with a €$10$ gift card.
@@ 4.4. Pilot User Study with Human Expert Demonstrators
5. Results and Discussion
5.1. Experiments with Oracle Experts
@@
5.2. Experiments with Human Experts
5.2.1. Learning Performance
@@
Figure 5. Distribution of provided demonstrations from one of the human participants using different baseline methods.